OpenStudy (anonymous):

Eliza started her savings account with $100. Each month she deposits $25 into her account. Write the function that describes the relationship between the amount Eliza deposits and the total amount in her account. Then, determine the average rate of change in the account from the 2nd month to the 10th month.

OpenStudy (anonymous):

Started with 100. (0, 100) Add 25: (1, 125) Add 25: (2, 150) Keep adding twenty-five till you get: (10, 350)

OpenStudy (anonymous):

Average rate of change formula: m=y2−y1x2−x1 y2=350 y1=150 x2=10 x1=2 m=350−15010−2 m=2008 so your answer is m=25

OpenStudy (anonymous):

If this helped can i have a medal? :D